A group of private investors purchased a condominium complex for $4 million. They made an initial down payment of 15%and obtained financing for the balance. The loan is to be amortized over 15 years at an interest rate of 12% per year compounded quarterly. (Round your answers to the nearest cent.)

What is the required quarterly payment?
$  

How much total interest will be paid on the loan?
$  
Answer #1

1)

Amount of loan = 4,000,000 (1 - 0.15)

Amount of loan = 3,400,000

Rate = 12%/ 4 = 3%

Number of periods = 15 * 4 = 60

Present value = Quarterly payment * [1 - 1 / (1 + r)n] / r

3,400,000 = Quarterly payment * [1 - 1 / (1 + 0.03)60] / 0.03

3,400,000 = Quarterly payment * [1 - 0.169733] / 0.03

3,400,000 = Quarterly payment * 27.675564

Quarterly payment = $122,852.06

2)

Total amount = 122,852.06 * 60 = 7,371,123.6

Total interest = 7,371,123.6 - 3,400,000

Total interest = $3,971,123.6